# ActErase: A Training-Free Paradigm for Precise Concept Erasure via Activation Patching

### v1

@@ B.2 Additional NSFW Erasure Results
For single concept erase, we have already presented the results of metrics in the main text. Here we give more generated images in Figure 9. We also evaluate the FID and clip score for target images or images generated by I2P dataset. FID is measured between images before and after erasing while clip score is calculated by erased images and prompts in I2P dataset. In theory, a lower clip score and higher FID illustrate lower similarity and better erasure performance. However, a higher FID value only indicates significant visual changes in the generated images but does not confirm that the semantics aligned with the target concept have been fully eliminated; conversely, this reduction may well indicate a degradation in image quality. Similarly, a lower CLIP score suggests that the generated images after erasure deviate substantially from the text prompts, yet it cannot specify whether this deviation corresponds to the intended removal of the target concept. In summary, while the FID and CLIP scores of images generated from the I2P dataset cannot directly measure erasure effectiveness, they remain valuable for further verification and comparison of the efficacy of different erasure methods.
Method FID ($\uparrow$) CLIP ($\downarrow$)
ESD 78.65 30.20
UCE 81.87 30.46
Concept-prune 81.31 30.43
AdaVD 82.14 30.95
STEREO 69.75 22.08
TRCE 82.09 27.56
[…]
For multiple concept erase, we have already shown the results of metrics in the main text. Here we show some generated images in figure 10. We also give the algorithm for multiple concept erasure in Algorithm 2. When erasing multi concept, we observe that compared to solely removing the ”nudity” concept, the model almost ceased generating human figures. We hypothesize that this occurs because the increase in the number of targeted concepts leads to a corresponding expansion in the affected activation parameters, thereby amplifying the impact on image generation. This ultimately results in the unintended and near-total erasure of the broader ”human” concept.
B.3 Additional Style Erasure Results
B.4 Additional Object Erasure Results
To evaluate object erase capabilities, we generated images for 10 concepts in ImageNet. For each erasing task, we generate 500 images for each concept with the prompt format ’a photo of ¡concept¿’, resulting in 500 images of the target concept and 4,500 images of other concepts. Here we will give details of each erasing task in Table 10. It illustrates that our method can erase different kinds of object concepts. We attached more generated images in Figure 14 and Figure 15.
@@ C Additional Analysis and Ablation Study
Ablation Study on Multiple Concept Erasure To further analyze the impact of the number of erased concepts on the performance of multiple concept erasure, we use 10,000 prompts from the MS-COCO dataset to generate images with multiple concept erasure models containing 1, 3, 5, and 10 object concepts respectively, and analyze the CLIP scores and FID metrics as shown in Table 11. As the number of concepts increases, the FID and CLIP Score decrease, and when the number of concepts reaches 10, the metrics indicate an almost complete failure to generate coherent images. We attribute this to the growing proportion of activated parameters introduced with more concepts, coupled with the distortion of these parameters during aggregation, which collectively disrupts the image generation process.
Number of concepts FID ($\uparrow$) CLIP ($\downarrow$)
1 20.85 30.72
3 43.86 28.78
5 84.37 25.54
10 222.14 20.34

### v2

@@ B.2 Additional NSFW Erasure Results
Algorithm 1 Concept Erasure via Activation Patching
Input: Pretrained model $G$, target concept $\mathbf{c}_{\text{target}}$, target prompt $c_{t}$, source prompt $c_{s}$, source concept $\mathbf{c}_{\text{source}}$, layer weight $\mathbf{W}_{s}$ and $\mathbf{W}_{t}$, layer set $\mathcal{L}$, timesteps $T$, threshold $\tau$
1:Step 1: Extract activations and generate masks.
[…]
3:for $t=T$ to 11 do
4:  $\mathbf{x}_{s}\leftarrow G(\mathbf{z}_{t},t,\mathbf{c}_{s})$
5:  $\mathbf{x}_{t}\leftarrow G(\mathbf{z}_{t},t,\mathbf{c}_{t})$
6:  for each layer $l\in\mathcal{L}$ do
8:   $I_{s}\leftarrow f(\mathbf{x}_{s},\mathbf{W}_{s})$
9:   $I_{t}\leftarrow f(\mathbf{x}_{t},\mathbf{W}_{t})$
10:   $\mathbf{M}^{l}\leftarrow(I_{s}\geq\tau)\land(I_{s}<I_{t})$
11:  end for
12:  $\mathbf{z}_{t-1}\leftarrow G(\mathbf{z}_{t},t,\mathbf{c}_{s})$
13:end for
14:return $\mathbf{x}_{\text{avg}},\ \mathbf{M}$
15:
16:Step 2: Apply patching.
[…]
18:for $t=T$ to 11 do
19:  $\mathbf{x}_{t}\leftarrow G(\mathbf{z}_{t},t,\mathbf{c}_{t})$
20:  for each layer $l\in\mathcal{L}$ do
22:  end for
24:end for
25:return $\mathbf{z}_{0}$
@@
For single concept erase, we have already presented the results of metrics in the main text. The algorithm of the activation patching method is shown in Algorithm 1. Here we give more generated images in Figure 10. We also evaluate the FID and clip score for target images or images generated by I2P dataset. FID is measured between images before and after erasing while clip score is calculated by erased images and prompts in I2P dataset. In theory, a lower clip score and higher FID illustrate lower similarity and better erasure performance. However, a higher FID value only indicates significant visual changes in the generated images but does not confirm that the semantics aligned with the target concept have been fully eliminated; conversely, this reduction may well indicate a degradation in image quality. Similarly, a lower CLIP score suggests that the generated images after erasure deviate substantially from the text prompts, yet it cannot specify whether this deviation corresponds to the intended removal of the target concept. In summary, while the FID and CLIP scores of images generated from the I2P dataset cannot directly measure erasure effectiveness, they remain valuable for further verification and comparison of the efficacy of different erasure methods.
@@
For multiple concept erase, we have already shown the results of metrics in the main text. Here we show some generated images in figure 11. We also give the algorithm for multiple concept erasure in Algorithm 2. When erasing multi concept, we observe that compared to solely removing the "nudity" concept, the model almost ceased generating human figures. We hypothesize that this occurs because the increase in the number of targeted concepts leads to a corresponding expansion in the affected activation parameters, thereby amplifying the impact on image generation. This ultimately results in the unintended and near-total erasure of the broader "human" concept.
B.3 Additional Style Erasure Results
B.4 Additional Object Erasure Results
To evaluate object erase capabilities, we generated images for 10 concepts in ImageNet. For each erasing task, we generate 500 images for each concept with the prompt format ’a photo of <concept>’, resulting in 500 images of the target concept and 4,500 images of other concepts. Here we will give details of each erasing task in Table 12. It illustrates that our method can erase different kinds of object concepts. We attached more generated images in Figure 15 and Figure 16.
@@ C Additional Analysis and Ablation Study
[…]
Table 14: Comparison of ’Nudity’ erasing results with different masks.
Type ASR(%) FID ($\uparrow$) CLIP ($\downarrow$) Without 0.0 84.57 25.75 Wanda 1.9 16.65 30.42 RAD(Relative Activation Difference) 16.59 27.27 25.93 MANU 2.5 15.75 30.36
@@
Ablation Study on type of masks To investigate the impact of mask types on the erasure effect of ActErase, we applied different mask generation methods under the same sparsity, as shown in Table 14 . It can be seen that masks generated simply using the Relative Activation Difference fail to erase the target concept, while masks obtained via the combined score in MANU and those derived from Wanda both achieve fairly good concept erasure. Under the same sparsity conditions, masks generated by the Wanda method exhibit the best erasure performance, whereas masks obtained through MANU’s combined score method demonstrate better generative capability.